A car initially traveling at 25.4 m/s undergoes a constant negative acceleration of magnitude 1.80 m/s2 after its brakes are applied.

(a) How many revolutions does each tire make before the car comes to a stop, assuming the car does not skid and the tires have radii of 0.330 m?
_______ rev

(b) What is the angular speed of the wheels when the car has traveled half the total distance?
________ rad/s
Answer #1

As

v^2 = vo^2 + 2ad

and vo = 25.4 m/s, a = -1.80 m/s^2, v = 0, solving for d,

d = 179.2 m

Thus, the number of revolutions is

N = d / [2 pi r]

Thus,

N = 86.4 revolutions   [ANSWER]

*********************

v^2 = vo^2 + 2ad

Then, if d = 179.2 m / 2 = 89.6 m,

v = 17.96 m/s

As

w = v / r

Then

w = 54.4 rad/s   [ANSWER]